OpenStudy (anonymous):

Does anyone know how to graph log3(x)? Medal will be given

OpenStudy (agent0smith):

\[\Large y = \log_3 x\]You can either use a table of values, and pick powers of 3 for x, eg. 1/9, 1/3, 1, 3, 9 and find y. Or use the change of base formula.

OpenStudy (anonymous):

ohh, ok thanks agent0smith

OpenStudy (anonymous):

me get it know

OpenStudy (mathmale):

First, let's review inverse functions: Although I won't go into much detail here, y =3^x and y=log(to the base 3) x are inverse functions, and thus cancel each other out. This property is useful in graphing y=log(to the base 3) (x). On the x-axis we'd want to mark powers of 3: 3^0, 3^1, 3^2, 3^3, and so on, or 1, 3, 9, 27, and so on. The corresponding values of y would then be y=log(to the base 3) (x), where we substitute 3^0, 3^1, 3^2, 3^3, and so on (which have the values 1, 3, 9, 27, and so on). So: if x=3^0, y=log(to the base 3) (3^0) = 0, or, alternatively, if x=1, y=0. If x=3^1, y=log(to the base 3) (3^1) = 1, or, alternatively, if x=3, y=1 If x=3^2, y=log(to the base 3) (3^2) = 2, or, alternatively, if x=9, y=2 If x=3^(-1), y=log(to the base 3) (3^(-1)) = -1, or, alternatively, if x=(1/3), y=-1

OpenStudy (mathmale):